OpenStudy (anonymous):

(4) to the power of 3 written out as -4x4x4x4

OpenStudy (anonymous):

-4x4x4

OpenStudy (anonymous):

ignore the the fourth four

OpenStudy (amistre64):

whats the - for?

OpenStudy (saifoo.khan):

why is there a negative sign for? :O

OpenStudy (amistre64):

but yes, that is what an exponent signifies

OpenStudy (anonymous):

-4

OpenStudy (amistre64):

4^3 = 4*4*4

OpenStudy (amistre64):

why would you want to include a spurious element into the picture?

OpenStudy (anonymous):

no becuase it is (-4) to the power of 3

OpenStudy (anonymous):

-4 ^3 =-64

OpenStudy (anonymous):

wait wat was this again ^ and how do you wirte that